OpenStudy (anonymous):

Match each form with its equation. 1. y = mx + b 2. Ax + By = C 3. x = c 4. y - y(with a little 1 right here) = m (x-x(and little 1) 5. y = b Answers: A. Slope- intercept form B. Point-slope form C. Standard form D. Horizontal E. Vertical Line

OpenStudy (anonymous):

it would be A Slope- intercept form

Nnesha (nnesha):

idk how i should explain without giving direct answer :P

OpenStudy (anonymous):

then just give the answer

jimthompson5910 (jim_thompson5910):

this is just something you have to memorize and know off the top of your head https://smsthinkingmaps.wordpress.com/2010/11/28/math-algebra-tree-map-forms-of-a-linear-equation-1-of-2/

Nnesha (nnesha):

oky for standard form all x and y should be on left side for slope intercept form see which is equation solved for standard form Ax+By=C slope intercept form y =mx+b point slope y-y1=m(x-x1)

Nnesha (nnesha):

for y***

OpenStudy (anonymous):

so i was right it would be a Yes

Nnesha (nnesha):

vertical line when x = something horizontal line when y = b hope you understand

OpenStudy (anonymous):

ok thank you

OpenStudy (anonymous):

Np

OpenStudy (anonymous):

you might want to look at there answer mine was wrong lol